1. (Chapter 6, question E12) A trait in garden peas involves the curling of leaves. A two-
factor cross was made by crossing a plant with yellow pods and curling leaves to a wild-type
plant with green pods and normal leaves. All the F1 offspring had green pods and normal leaves.
The F1 plants were then crossed to plants with yellow pods and curling leaves. The following
results were obtained:
117 green pods, normal leaves
115 yellow pods, curling leaves
78 green pods, curling leaves
80 yellow pods, normal leaves
A. Conduct a Chi-Square analysis to determine if these two genes are linked.
B. If they are linked, calculate the map distance between the two gene
Answer #1

Recombinant frequency(RF) =78+80/390

=158/390

= 0.40

This seems like a case of linkage because the RF is less than 50% expected from independent assortment.

Total progeny is 390 , we would expect 390/4 = 97 in each class

A.) Chi square = 3.3+4.12+3.721+2.97

= 14.11

Degree of freedom(DF) = 4-1=3

In table at P = 0.05 and DF = 3, value = 7.815

Since our chi square value is greater than the table value genes are linked.

B.) Map distance = percentage recombination

= 78+80/390x100

=  40 mu